Expected Value and Fair Price. Your friend wants to sell you a raffle ticket she bought for 10 dollars. It is possible to win one of two prizes in the raffle: you can either win $200, or win $50 (or win nothing). 1 out of 100 (1%; .01) entrants will win $200, while 10 more people out of 100 (10%; .1) will win $50. a) Find the expected value of paying your friend 10 dollars for the raffle ticket (let B=buying the $10 ticket). Exp(B)=____________________________________________________________________

b) Find the fair price of the ticket:______________________________________________
Answer #1

here expected value =E(B) =expeced win -cost of ticket =0.01*200+0.1*50-10= -3

b)fair price =expected win =0.01*200+0.1*50 =7